OpenStudy (swag):

100 plus the product of a number and -2 equals 50. What is the number?

OpenStudy (theviper):

Product of a no. ? How is this possible ?

OpenStudy (swag):

Not sure

OpenStudy (anonymous):

100+(-2x) = 50 -50=-2x x=25

OpenStudy (theviper):

@MFace how can u say that the no. is multiplied by 2 ?

OpenStudy (anonymous):

the product is of a number (x) and -2

OpenStudy (swag):

o.o